OpenStudy (anonymous):

(2x+4y)/(x^2-y^2) times (x+y)/(x^2+4xy+4y^2) just checking is this the right answer? 2/(x62-y^2) ?

OpenStudy (anonymous):

sorry I meant 2/(x^2-y^2)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok could you explain what I got wrong?

OpenStudy (anonymous):

\[2/(x^2 +xy-2y^2)\]

OpenStudy (anonymous):

I need to see your solution in order to find your mistake/ It's just regular stuff

OpenStudy (anonymous):

Here's how I factored it out: \[2(x+2y)/(x+y)(x-y) \times (x+y)/(x+2y)(x+2y)\]

OpenStudy (anonymous):

then, I removed (x+y) from denominator and numerator, and then (x+2y) to simplify

OpenStudy (anonymous):

leaving me with 2/(x-y)(x+2y)

OpenStudy (anonymous):

then I multiplied the fractions

OpenStudy (anonymous):

ohhh wait now I see what I got wrong...I forgot about the 2 factor in 2y...but thanks anyway... :)

OpenStudy (anonymous):

u got mistake at the very end? welcome

OpenStudy (anonymous):

\[(x-y)(x+2y)=x^2+xy-2y^2\]

OpenStudy (anonymous):

(x^2-y^2) is equal to (x+y)(x-y), not (x-y)(x+2y)

OpenStudy (anonymous):

sorry, i had to sign up to reply so it took super long lol

OpenStudy (anonymous):

its ok lol thanks for trying! haha